# references.bib

### v1

@@ 3 Method
Regression objective. To realize the downstream regression task, i.e., regressing parameters of an underlying governing equation for collective behavior, we have multiple choices. By our assumption of a latent dynamic process that carries information about the dynamic nature of the topological changes, it is reasonable to tie the regression objective to the latent path. In particular, we can even re-use the attention-based encoder architecture to allow attending to different parts of this path. However, different to $\text{{Enc}}_{\boldsymbol{\theta}}$, which parametrizes the approximate posterior from observations at time points $\tau_{i}$, the regression network $\text{{Enc}}_{\boldsymbol{\alpha}}$ (with its own set of parameters $\boldsymbol{\alpha}$) receives latent states at equidistant $t_{i}\in[0,T]$. As in [Rubanova19a], for training, we extend the ELBO objective of Eq. 4 by an additive auxiliary regression loss, such as the mean-squared error.
A schematic overview of our neural persistence dynamics approach is shown in Fig. 1; additional details, including different model variants, are illustrated in Fig. 2.
@@ 4.1 Datasets
Similar to previous work [Topaz15a, Bhaskar19a, Xian22a, Giusti23a], we simulate models of collective behavior under different parametrizations to produce a large corpus of observation sequences with varying dynamics and varying emergent behavior. For simulation, we use SiSyPHE [Diez21a] [^2] https://github.com/antoinediez/Sisyphe and consider three different models in $\mathbb{R}^{3}$: D’Orsogna [Dorsogna06a], Vicsek [Degond08a] and volume exclusion [Motsch18a]. The corresponding equations of motion (and interaction laws) are summarized in Fig. 3 with parameters that are varied in our experiments highlighted in red. We sample these parameters, as specified in Appendix A, so that a wide range of macroscopic behavioral regimes is covered. For each sampled parameter configuration, we simulate one sequence of point clouds, to a total of 10,000 sequences per model. All simulations are run for 1,000 time steps (with step size 0.01, starting at $t=0$) on point clouds of size $M=200$. We take every 10th time step as an observation, yielding observation sequences of length 100. At $t=0$, points are drawn uniformly in $[-0.5,0.5]^{3}$, and initial velocities are uniformly distributed on the unit sphere. For direct comparison to [Giusti23a], we also simulated their parameter configuration of the D’Orsogna model. This setup yields four datasets: dorsogna-1k (from [Giusti23a]), dorsogna-10k, vicsek-10k and volex-10k.
@@ 4.2 Implementation, training & evaluation metrics
In general, we randomly partition each dataset into five 80/20 splits for training and testing. On each training split, we train three models using only a fraction (i.e., 20%, 50% and 80%) of randomly chosen time points per sequence. Similarly, all testing splits undergo the same sampling procedure. Hence, our evaluation metrics (see below) provide a robust estimate of performance for different regimes of missing and unevenly spaced observations. All models are trained using ADAM [Kingma15a] with a weight decay and KL divergence weight of 0.001 for 150 epochs at a learning rate of 0.001. To compute Vietoris-Rips persistence diagrams for zero, one, and two-dimensional persistent homology, we rely on Ripser++ [Zhang20a]. Vectorizations for each diagram dimension are computed using exponential structure elements from [Hofer19b, Sec. 5.4.1] (i.e., the $s_{\boldsymbol{\lambda}}$ in Eq. 3), located at 20 $k$-means++ cluster centers and with variance(s) chosen according to [Royer21a]. Upon concatenation of the vectorizations, this yields $d=60$ dimensional representations per point cloud and time point. Latent space dimensionality $z$ is fixed to 20 and the ODE integration time is always scaled to $[0,1]$. While other settings are certainly possible, we did not observe any noticeable benefits from increasing the dimensionality of the vectorizations or the latent space.
Evaluation metrics. We report the coefficient of determination $R^{2}$ (in $[0,1]$; higher is better $\uparrow$) and the symmetric mean absolute percentage error (SMAPE, in $[0,1]$; lower is better $\downarrow$). In detail, we first collect both scores on the testing portion of our five random 80/20 splits, and all three time point sampling percentages. In each table, we then list the average of these scores with one standard deviation. Further, when reporting results on a dataset, we mark the best score in bold, as well as all other scores that do not show a statistically significant difference in mean (assessed via a Mann-Whitney test at 5% significance, and correcting for multiple comparisons).
@@ 4.3 Ablation
In our ablation study, we assess (1) the relevance of different point cloud representations (PH vs. PointNet++), (2) any potential benefits of modeling latent dynamics and (3) the impact of varying the observation timeframe.
@@
Source $\oslash\ R^{2}\uparrow$oslash SMAPE $\downarrow$
@@
PointNet++ (v2) 0.814$\pm$0.032 0.132$\pm$0.018
PH (v1) 0.846$\pm$0.011 0.097$\pm$0.005
PH+PointNet++ (v3) 0.930$\pm$0.003 0.067$\pm$0.002
Tbl. 2 presents an extension of Tbl. 1, listing results for the vicsek-10k and dorsogna-1k data. While using PointNet++ representations on vicsek-10k yields rather poor performance, combining them with representations computed from PH is at least not detrimental. On the other hand, on dorsogna-1k, where PointNet++ yields decent SMAPE and $R^{2}$ scores, the combination of both sources substantially outperforms each single source in isolation. Although, the complementary nature of a topological perspective on data has been pointed out many times in the literature, it is rarely as pronounced as in this particular experiment. Hence, we will stick to the combination of PH+PointNet++ representations (i.e., v3 in Fig. 2) in any subsequent experiments.
Are explicit latent dynamics beneficial? To assess the impact of modeling the dynamics of persistence diagram vectorizations, we ablate the latent ODE part of our model. In detail, we compare v3 from Fig. 2 against using $\texttt{Enc}_{\boldsymbol{\alpha}}$ operating directly on concatenated point cloud representations from PH and PointNet++ (i.e., a combination of v4 & v5 from Fig. 2). The latter approach constitutes an already strong baseline (cf. [Zeng23a]) as $\texttt{Enc}_{\boldsymbol{\alpha}}$ incorporates an attention mechanism that allows attending to relevant parts of each sequence. For a fair comparison, we increase the size of $\texttt{Enc}_{\boldsymbol{\alpha}}$ to approximately match the number of parameters to our latent dynamic model.
$\oslash\ R^{2}\uparrow$oslash SMAPE  $\downarrow$
@@
w/ dynamics 0.930$\pm$0.003 0.067$\pm$0.002
w/o dynamics 0.850$\pm$0.006 0.098$\pm$0.003
Table 3: Results on the relevance of latent dynamics.
Similar to Tbl. 2, Tbl. 3 lists results for the vicsek-10k and dorsogna-1k data. As we can see, explicitly modeling the latent dynamics is beneficial, with considerable improvements in the reported scores across both datasets. While differences are quite prominent on the dorsogna-1k data, they are less pronounced in the SMAPE score on vicsek-10k, but still statistically significant. However, it is important to point out, that even without any explicit latent dynamics, the regression performance is already above the current state-of-the-art (cf. PSK [Giusti23a] results in Tbl. 4), highlighting the necessity for strong baselines.
Impact of the observation timeframe. So far, we have presented results (as in prior work) where observation sequences are extracted from the beginning of a simulation (i.e., from $t_{0}=0$), which corresponds to observing the emergence of patterns out of a random, yet qualitatively similar, initial configuration of points. In this experiment, we investigate the impact of extracting observation sequences with starting points randomly spread across a much longer simulation time. This is relevant as it is typically unlikely to observe a random configuration of points in practice initially. To this end, we create variations of the dorsogna-10k data by progressively increasing the simulation time $T$ from 1,000 to 20,000, and then randomly extracting sub-sequences of length 1,000 (again, as before, taking each 10th step as an observation). The extension of the simulation timeframe has the effect that the difficulty of the learning task considerably increases with $T$. We argue that this is due to the increased variation across the observed sequences, while the amount of training data remains the same. For comparison, we also list results for the PSK approach of [Giusti23a], however, only for the optimal case of observing all time points within a sequence (due to the massive PSK computation time; >3days). As the PSK relies on persistent homology only, we use our v1 variant from Fig. 2. Fig. 4 displays an initial increase/drop in SMAPE and $R^{2}$, resp., (cf. Tbl. 4) for both approaches, yet our approach degrades much slower with $T$.
2,0005,00010,00020,$0$.40.40.40.40.50.50.50.50.60.60.60.6Simulation time $T$$R^{2}~{}(\uparrow)$Ours, v1PSK [Giusti23a]
@@
Figure 4: Impact of the maximal simulation time $T$ for extracting training/testing sequences starting at $\tau_{0}\in[0,T-1000]$.
@@ 4.4 Comparison to the state-of-the-art
Finally, we present parameter regression results on the full battery of datasets and compare against two of the most common approaches from the literature: the path signature kernel (PSK) of [Giusti23a] and the crocker stacks approach from [Xian22a]. For both approaches, we report results when all time points are observed to establish a baseline of what could be achieved in the best case. Note that we evaluate the PSK approach [Giusti23a] on exactly the same vectorizations as our approach and we replicate their experimental setting of choosing the best hyperparameters via cross-validation. Similarly, for crocker stacks, we cross-validate the discretization parameters, see Secs. B.1 and B.2. Tbl. 4 lists the corresponding results. Aside from the observation that our approach largely outperforms the state-of-the-art on these parameter regression tasks, we also highlight that our model is trained with exactly the same hyperparameters across all datasets.
@@
$\oslash\,R^{2}\uparrow$oslash SMAPE $\downarrow$
dorsogna-1k Ours (joint, v3) 0.930$\pm$0.003 0.067$\pm$0.002
Ours (PH-only, v1) 0.846$\pm$0.011 0.097$\pm$0.005
PSK [Giusti23a] 0.816$\pm$0.015 0.096$\pm$0.006
Crocker Stacks [Xian22a] 0.743$\pm$0.083 0.067$\pm$0.044
@@
PSK [Giusti23a] 0.641$\pm$0.006 0.100$\pm$0.003
Crocker Stacks [Xian22a] 0.258$\pm$0.316 0.294$\pm$0.405
vicsek-10k Ours (joint, v3) 0.574$\pm$0.030 0.144$\pm$0.006
Ours (PH-only, v1) 0.578$\pm$0.032 0.146$\pm$0.006
@@
Crocker Stacks [Xian22a] 0.342$\pm$0.167 0.878$\pm$0.519
@@
Crocker Stacks [Xian22a] -0.359$\pm$0.842 0.177$\pm$0.074
@@
A closer look at the volex-10k dataset, in particular, its governing equations in Fig. 3, shows that the cardinality of the point clouds may change over time due to cell division or death. While this is no practical limitation for our approach, our stability arguments from Eq. 3 no longer apply. We conjecture that one may be able to extend the result of [Skraba23a] to account for such cases, but this might require a case-by-case analysis and possible include additional assumptions.
@@ 5 Discussion
We introduced a framework to capture the dynamics observed in topological summary representations over time and presented one incarnation using a latent ODE on top of vectorized Vietoris-Rips persistence diagrams, addressing the problem of predicting parametrizations for models of collective behavior. Our approach successfully scales to a large number of observation sequences, requires little to no parameter tuning, and largely outperforms the current state-of-the-art, as demonstrated by several experiments on dynamic point cloud datasets with varying characteristics.
Limitation(s). One obvious limitation in the presented implementation is the reliance on Vietoris-Rips persistent homology of point clouds. In fact, the underlying simplicial complex will become prohibitively large (especially for $H_{2}$) once we scale up the number of points by, e.g., an order of magnitude. Using an Alpha [Edelsbrunner2010Computational] or a Witness complex [DeSilva04a] might be a viable alternative to mitigate this issue. Similarly, one may explore subsampling strategies for persistent homology, as in [Chazal15a], and learn a latent ODE from a combination of estimates.
Finally, we would like to reiterate that the core idea of learning the dynamics of vectorized persistence diagrams via a continuous latent variable model is not tied to one specific problem setting. In fact, one may apply our approach to other types of time-varying data (e.g., graphs) with appropriate downstream objectives. We hope that our work will inspire further research in this direction.

### v2

@@ 3 Method
Regression objective. To realize our downstream regression task, i.e., predicting parameters $\beta_{1},\dots,\beta_{P}$ of an underlying governing equation for collective behavior (see Fig. 3) from a given observation sequence, we have multiple choices. By our assumption of a latent dynamic process that carries information about the dynamic nature of the topological changes, it is reasonable to tie the simulation parameter estimates $\hat{\beta}_{1},\dots,\hat{\beta}_{P}$ to the latent path $\{\mathbf{z}_{t_{i}}\}_{i}$ via a regression network $\texttt{Reg}_{\boldsymbol{\alpha}}$ that accepts $\{\mathbf{z}_{t_{i}}\}_{i}$ as input. In particular, we re-use the attention-based encoder architecture Encθ to allow attending to different parts of this path. However, different to Encθ, which parametrizes the approximate posterior from observations at time points $\tau_{i}$, the regression network $\text{{Reg}}_{\boldsymbol{\alpha}}$ (with its own set of parameters $\boldsymbol{\alpha}$) receives latent states $\mathbf{z}_{t_{i}}$ at equidistant $t_{i}\in[0,T]$, then summarizes this sequence into a vector and linearly maps the latter to $\hat{\beta}_{1},\dots,\hat{\beta}_{P}$, cf. Fig. 2. As in [Rubanova19a], for training, we extend the ELBO objective of Eq. 4 by an additive auxiliary regression loss, such as the mean-squared error (MSE), between the predictions $\hat{\beta}_{p}$ and the ground truth $\beta_{p}$, implicitly making a Gaussian noise assumption.
A schematic overview of our Neural Persistence Dynamics approach is shown in Fig. 1. Additional details, including different model variants, are illustrated in Fig. 2.
Remark 2.
Clearly, our presented framework allows for many architectural choices. While some components affect downstream (regression) performance only marginally, others have a more profound impact, and we have already identified some recommended choices above (based on our experiments in Sec. 4). This includes (1) our choice of a stable persistent homology vectorization $\operatorname{vec}$, where we rely on [Hofer19b] due to consistently reliable performance without much hyperparameter tuning, and (2) our choice of recognition network, where we choose an attention-based approach (mTAN) [Shukla21a] which has proven to be very effective in practice [Zeng23a]. In Sec. 4.2, we will provide additional configuration details for our experimental study.
@@ 4.1 Datasets
Similar to previous work [Topaz15a, Bhaskar19a, Xian22a, Giusti23a], we evaluate and compare our approach on simulation data that is generated from parametric models of collective behavior. Specifically, we consider three different models in $\mathbb{R}^{3}$: D’Orsogna [Dorsogna06a], Vicsek [Degond08a] and volume exclusion [Motsch18a], using the publicly-available implementations in the SiSyPHE library [Diez21a] [^2] https://github.com/antoinediez/Sisyphe. The corresponding equations of motion (and interaction laws) are summarized in Fig. 3. The parameters that are varied to generate the datasets, i.e., the response variables of the regression tasks, are highlighted in red. We sample these parameters, as specified in Appendix A, to cover a wide range of macroscopic behavioral regimes. For each sampled parameter configuration, we simulate one sequence of point clouds, to a total of 10,000 sequences per model. All simulations are run for 1,000 time steps (with step size 0.01, starting at $t=0$) on point clouds of size $M=200$. We take every 10th time step as an observation, yielding observation sequences of length 100. At $t=0$, points are drawn independently and uniformly in $[-0.5,0.5]^{3}$, and initial velocities are uniformly distributed on the unit sphere. For direct comparison to [Giusti23a], we also simulated their parameter configuration of the D’Orsogna model. This setup yields four datasets: dorsogna-1k (from [Giusti23a]), dorsogna-10k, vicsek-10k and volex-10k.
@@ 4.2 Implementation, training & evaluation metrics
Implementation. The model variants from Fig. 2 can be realized in many ways. Below, we specify the configuration that was used to run the experiments. For the encoder $\text{{Enc}}_{\boldsymbol{\theta}}$, as well as the regression network $\text{{Reg}}_{\boldsymbol{\alpha}}$, we use the attention-based mTAN architecture [Shukla21a]. As decoder network $\text{{Dec}}_{\boldsymbol{\gamma}}$, we choose a two-layer MLP with ReLU activations, and as ODE solver, we select the Euler method. Each model is trained for 150 epochs using ADAM [Kingma15a] (with a weight decay of 0.001), starting at a learning rate of 0.001 (decaying according to a cosine annealing schedule) and MSE as a reconstruction (i.e., to evaluate the first term in Eq. 4) and regression loss. In case a model uses topological features, we use Ripser++ [Zhang20a] to compute (prior to training) persistent homology of dimension up to two, i.e., $\operatorname{H}_{0}$, $\operatorname{H}_{1}$, and $\operatorname{H}_{2}$. Vectorizations of each persistence diagram are then obtained using exponential structure elements from [Hofer19b, Def. 19]. In particular, we use 20 structure elements per diagram, which yields a $d=3\!\cdot\!20$ dimensional representation per point cloud and time point. The location of each structure element is set to one of 20 $k$-means++ cluster centers, obtained by running the latter on a random subset of 50,000 points (per dimension) selected from all persistence diagrams available during training; the scale parameter of each structure element is set according to [Royer21a, Eq. (2)]. To model the dynamics, we fix the latent space dimensionality to $z=20$ and scale the ODE integration time to $[0,1]$. While other settings are undoubtedly possible, we did not observe any noticeable benefits from increasing the dimensionality of the vectorizations or the latent space. Our publicly available reference implementation can be found at https://github.com/plus-rkwitt/neural_persistence_dynamics.
Evaluation metrics. We randomly partition each dataset into five training/testing splits of size 80/20. To obtain a robust performance estimate for different regimes of missing and unevenly spaced observations, we train three models (per split) using only a fraction (i.e., 20%, 50%, and 80%) of randomly chosen time points per sequence. Similarly, all testing splits undergo the same sampling procedure. All scores (see below) are reported as an average ($\pm$ one standard deviation) over the five splits and the three time point sampling percentages. Specifically, we report the variance explained (VE) [Li17a] (in $[0,1]$; higher is better $\uparrow$) and the symmetric mean absolute percentage error (SMAPE) (in $[0,1]$; lower is better $\downarrow$). For each testing sequence in each split, the scores are computed from the true simulation parameters (see variables marked red in Fig. 3) and the corresponding predictions (denoted by $\hat{\beta}_{i}$ in Fig. 2). Finally, when reporting results on a dataset, we mark the best score in bold, as well as all other scores that do not show a statistically significant difference in mean (assessed via a Mann-Whitney test at 5% significance and correcting for multiple comparisons).
@@ 4.3 Ablation
In our ablation study, we assess (1) the relevance of different point cloud representations (PH vs. PointNet++), (2) any potential benefits of modeling latent dynamics and (3) the impact of varying the observation timeframe. Additional ablation experiments can be found in Sec. B.3.
@@
[…]
@@
PointNet++ (v2) 0.816$\pm$0.031 0.132$\pm$0.018
PH (v1) 0.851$\pm$0.008 0.097$\pm$0.005
PH+PointNet++ (v3) 0.931$\pm$0.004 0.067$\pm$0.002
Tbl. 2 shows an extension of Tbl. 1, listing results for the vicsek-10k and dorsogna-1k data. While using PointNet++ representations on vicsek-10k yields rather poor performance, combining them with representations computed from PH is at least not detrimental. On the other hand, on dorsogna-1k, where PointNet++ yields decent SMAPE and VE scores, the combination of both sources substantially outperforms each single source in isolation. Although, the complementary nature of a topological perspective on data has been pointed out many times in the literature, it is rarely as pronounced as in this particular experiment. Hence, we will stick to the combination of PH+PointNet++ representations (i.e., v3 in Fig. 2) in any subsequent experiments.
Are explicit latent dynamics beneficial? To assess the impact of explicitly modeling the dynamics of persistence diagram vectorizations, we ablate the latent ODE part of our model. In detail, we compare v3 from Fig. 2 against using $\texttt{Reg}_{\boldsymbol{\alpha}}$ operating directly on concatenated point cloud representations from PH and PointNet++ (i.e., a combination of variants v4 & v5 from Fig. 2). The latter approach constitutes an already strong baseline (cf. [Zeng23a]) as Regα incorporates an attention mechanism that allows attending to relevant parts of each sequence. For a fair comparison, we increase the size of $\texttt{Reg}_{\boldsymbol{\alpha}}$ to approximately match the number of parameters to our latent dynamic model.
Table 3: Results on the relevance of latent dynamics.
[…]
@@
w/ dynamics 0.931$\pm$0.004 0.067$\pm$0.002
w/o dynamics 0.850$\pm$0.006 0.098$\pm$0.004
As in Tbl. 2, we list results for the vicsek-10k and dorsogna-1k data in Tbl. 3. First, we see that explicitly modeling the latent dynamics is beneficial, with considerable improvements in the reported scores across both datasets. While differences are quite prominent on the dorsogna-1k data, they are less pronounced in the SMAPE score on vicsek-10k, but still statistically significant. Second, it is important to point out that even without any explicit latent dynamics, the regression performance (see Tbl. 4) is already above the current state-of-the-art (i.e., compared to PSK [Giusti23a]), highlighting the necessity for strong baselines.
Impact of the observation timeframe. So far, we have presented results (as in prior work) where observation sequences are extracted from the beginning of a simulation (i.e., starting from $t_{0}=0$), which corresponds to observing the emergence of patterns out of a random yet qualitatively similar, initial configuration of points. In the following experiment, we investigate the impact of extracting observation sequences with starting points randomly spread across a much longer simulation time. This is relevant as it is unlikely to observe uniform initial positions in any data obtained from real-world experiments. To this end, we create variations of the dorsogna-10k data by progressively increasing the simulation time $T$ from 1,000 to 20,000 and then randomly extracting sub-sequences of length 1,000 (again, as before, taking each 10th step as an observation). The extension of the simulation timeframe has the effect that the difficulty of the learning task considerably increases with $T$. We argue that this is due to the increased variation across the observed sequences while the amount of training data remains the same. For comparison, we also list results for the PSK approach of [Giusti23a], however, only for the optimal case of observing all time points within a sequence (due to the massive PSK computation time; $>$ 3 days). As the PSK relies on persistent homology only, we limit our model to the v1 variant from Fig. 2. As seen from Fig. 4, we observe an increase/drop in SMAPE and VE, resp., for both approaches, yet our approach degrades much slower with $T$.
2,0005,00010,00020,$0$.40.40.40.40.50.50.50.50.60.60.60.6Simulation time $T$VE ()($\uparrow$)( ↑ )Ours, v1PSK [Giusti23a]
@@
Figure 4: Impact of the maximal simulation time $T$ for extracting training/testing sequences starting at $\tau_{0}\in[0,T-1000]$, assessed on the dorsogna-10k dataset.
@@ 4.4 Comparison to the state-of-the-art
Finally, we present parameter regression results on the full battery of datasets and compare against two of the most common approaches from the literature: the path signature kernel (PSK) of [Giusti23a] and the crocker stacks approach from [Xian22a]. For both competitors, we report results when all time points are observed to establish a baseline of what could be achieved in the best case. Note that we evaluate the PSK approach [Giusti23a] on exactly the same vectorizations as our approach, and we replicate their experimental setting of choosing the best hyperparameters via cross-validation. Similarly, for crocker stacks, we cross-validate the discretization parameters, see Secs. B.1 and B.2. Tbl. 4 lists the corresponding results. Aside from the observation that our approach largely outperforms the state-of-the-art across all parameter regression tasks, we also highlight that our model is trained with exactly the same hyperparameters on all datasets.
@@
[…]
dorsogna-1k Ours (joint, v3) 0.931$\pm$0.004 0.067$\pm$0.002
Ours (PH-only, v1) 0.851$\pm$0.008 0.097$\pm$0.005
PSK [Giusti23a] 0.828$\pm$0.016 0.096$\pm$0.006
Crocker Stacks [Xian22a] 0.746$\pm$0.023 0.150$\pm$0.005
@@
PSK [Giusti23a] 0.647$\pm$0.005 0.100$\pm$0.003
Crocker Stacks [Xian22a] 0.343$\pm$0.016 0.145$\pm$0.001
vicsek-10k Ours (joint, v3) 0.576$\pm$0.030 0.144$\pm$0.006
Ours (PH-only, v1) 0.579$\pm$0.034 0.146$\pm$0.006
@@
Crocker Stacks [Xian22a] 0.345$\pm$0.005 0.190$\pm$0.001
@@
Crocker Stacks [Xian22a] 0.076$\pm$0.019 0.292$\pm$0.004
@@
A closer look at the volex-10k dataset, in particular its governing equations in Fig. 3, shows that the cardinality of the point clouds may change over time due to cell division or death. While this is no practical limitation for our approach, our stability arguments from Eq. 3 no longer apply. We conjecture that one may be able to extend the result of [Skraba23a] to account for such cases, but this might require a case-by-case analysis and possibly include additional assumptions.
@@ 5 Discussion
We introduced a framework to capture the dynamics observed in topological summary representations over time and presented one incarnation using a latent ODE on top of vectorized Vietoris-Rips persistence diagrams, addressing the problem of predicting parametrizations for models of collective behavior. Conceptually, Neural Persistence Dynamics embodies the idea of learning the dynamics in a temporal sequence of vectorized topological summaries instead of, e.g., trying to track individual homology classes over time (as with persistence vineyards, see Sec. 2). Our approach successfully scales to a large number of observation sequences, requires little to no parameter tuning, and vastly outperforms the current state-of-the-art, as demonstrated by several experiments on dynamic point cloud datasets with varying characteristics. Finally, we emphasize that the fundamental ideas of Neural Persistence Dynamics may also be applied to other types of time-varying data (e.g., graphs or images) and downstream objectives simply by adjusting the filtration choice. We hope that our work will stimulate further research in this direction.
Limitation(s). One obvious limitation in the presented implementation is the reliance on Vietoris-Rips persistent homology of point clouds. In fact, the underlying simplicial complex will become prohibitively large (especially for $\operatorname{H}_{2}$) once we scale up the number of points by, e.g., an order of magnitude. Using an Alpha [Edelsbrunner2010Computational] or a Witness complex [DeSilva04a] might be a viable alternative to mitigate this issue. Similarly, one may explore subsampling strategies for persistent homology, as in [Chazal15a], and learn a latent ODE from a combination of estimates.
@@
Acknowledgments
This work was supported by the Land Salzburg within the EXDIGIT project 20204-WISS/263/6-6022 and projects 0102-F1901166- KZP, 20204-WISS/225/197-2019. M. Uray and S. Huber are supported by the Christian Doppler Research Association (JRC ISIA). We also thank all reviewers for the valuable feedback during the review process.